Isaac applies a force of 40 N at an angle of 60 degrees up from the horizontal to a wooden rod using a spring scale. If he generates a torque of 73m.N, how long was the rod? Show the givens, required, equations to be used, calculations and final answer

1
Expert's answer
2021-11-18T12:21:47-0500

The torque is


"T=Fd\sin\theta,"

express the length:

"d=\frac{T}{F\sin\theta}=2.1\text{ m}."